OpenStudy (anonymous):

u substitution, anti derivative of y/(y+1)^(3) dy

OpenStudy (freckles):

u=y+1 u-1=y du=dy

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{u-1}{u^3} du=\int\limits_{}^{}(u^{-2}-u^{-3}) du\]

OpenStudy (anonymous):

nevermind i got it